the physician has ordered 75 ml a 1:30 solution of acetic acid. How would you prepare this solution from a 5% solution of acetic acid? explain in a clear sentence
Answer by ewatrrr(24785)


Hi
1:30 solution of acetic acid is a .0333% solution or .9777% water
One would need to dilute the proper amount of 5% solution to achieve this.
Let x be the amount of the 5% solution to be used(95% water)
.95x + (75-x) = .9777*75ml
75-.9777*75 = .05x
33.45ml = x, the amount
33.45ml of 5% solution would needed to be mixed with 41.55ml of water.
.95*33.45 + 41.55 = 31.7775 + 41.55 = .9777*75 = 73.3275